OpenStudy (anonymous):

Find the probability of correctly answering the first 2 questions on a multiple choice test if random guesses are made and each question has 5 possible answers.

OpenStudy (anonymous):

\[\frac{1}{5}\times \frac{1}{5}\]

OpenStudy (anonymous):

5 possibilities 1 good outcome means 1/5 chance of getting 1 question right As the first and the second question are independent events there chance of happening together is their multiple: 1/5 *1/5=1/25

OpenStudy (anonymous):

thanks guys!